Which graph represents the parametric equations x = 3t and t3 – 1, where –2 ≤ t ≤ 3?



1
Expert's answer
2022-02-13T17:47:34-0500

"x=3t, y=t^3-1, t\in[-2,3]\\
x\in[-6,9],y\in[-9,26]\\

t=\frac{x}{3}\\
y=(\frac{x}{3})^3-1=\frac{x^3}{27}-1"

"y=\frac{x^3}{27}-1\\
\begin{matrix}
 x &-6&-3&0&3&6&9 \\
 y & -9&-2&-1&0&9&26
\end{matrix}"